Pick out the solution from the values given in the bracket next to each equation. Show that the other values do not satisfy the equation.
q / 2 = 7                        (7, 2, 10, 14)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

q / 2 = 7
q = 14 is a solution for this equation because for q = 14
q / 2 = 14 / 2
∴Equation satisfied
q = 7 is not a solution for this equation because for q = 7,
q / 2 = 7 / 2 and not 7
q = 2 is not a solution for this equation because for q = 2,
q / 2 = 2/ = 1 and not 7
q = 10 is not a solution for this equation because for q = 10,
q / 2 = 10 / 2
= 5 and not 7

Thanks